If mass of earth is same and the radius is decreased by 1%,then what will we the' g' at earth temprature
 

Grade:10

1 Answers

Khimraj
5 years ago
accelerationdue to gravity is given by
g = GM/R2
\Deltag/g = \DeltaM/M + 2\DeltaR/R
then % change in g = 0 + 2*1 = 2%
Hope it clears. If you like answer then please approve it.